media%2Ff6b%2Ff6b4d3ee-9cf7-4115-8604-a8
Answer #1

k\: =\: \frac{2.303}{t}\; log\: \frac{P_{0}}{(2P_{0}-P_{t})}

k\: =\: \frac{2.303}{12*60}\; log\: \frac{0.850}{(2*0.850-0.650)}

k\: =\: \frac{2.303}{720}\; log\: \frac{0.850}{(1.7-0.650)}

k\: =\: 0.003198\; log\: \frac{0.850}{(1.7-1.05)}

k\: =\: 0.003198\; log\: \frac{0.850}{(0.65)}

k\: =\: 0.003198\; log\: 1.3076

k\: =\: 0.003198\; *\: 0.1164

k\: =\: 0.0003722\: \; sec^{-1}

ii) Half life period

t_{\frac{1}{2}}\; =\: \frac{0.693}{k}

t_{\frac{1}{2}}\; =\: \frac{0.693}{0.0003722}

t_{\frac{1}{2}}\; =\: 1861.90\: sec^{-1}